OpenStudy (anonymous):

How many grams of H2O are necessary to weigh out 1 mole of H2O?

OpenStudy (rushwr):

okai ! SO fist of all u have to know this equation \[n= \frac{ m }{ M }\] Where n is no. of moles m is mass M is molar mass We know no. of moles, that is one mole and we have to find the molar mass which is 18g per mole So now you have to substitute the values and get the answer :)

OpenStudy (rushwr):

If \[n = \frac{ m }{ M }\] then m= n *M so m = 1mole * 18gmol^-1 = 18g

OpenStudy (rushwr):

Did you get it ?

OpenStudy (michele_laino):

hint: the mass of 1 molecule of \(H_2O\) is \(18\) \(amu\)